OpenStudy (anonymous):

identify the solution of the system of equations if any 3x-2y=6 5x=5-25y

OpenStudy (texaschic101):

5x = 5 - 25y -- divide both sides by 5 x = 1 - 5y now sub 1 - 5y in for x in the other equation to find y. Can you do this ?

OpenStudy (texaschic101):

let me start you off... 3x - 2y = 6 3(1 - 5y) - 2y = 6 -- now distribute through the parenthesis and then solve for y. Once you have y, sub that back into either equation to find x.

OpenStudy (anonymous):

is y -6